# Monitoring-feedback induced entanglement relaxations in a tilted free fermionic chain

### v1

@@ Abstract
Recent years have witnessed measurement-induced entanglement phase transitions attracting significant attention, yet the feedback-induced skin effect in tilted fields remains understudied. This work investigates entanglement relaxation and dynamical transitions in a fermionic chain subject to generalized monitoring and an applied tilted field. We show that the feedback-induced skin effect persists with tilting, though relaxation time increases. For small tilts, bipartite entanglement exhibits a widened metastable phase, transitioning from a logarithmic-law to an area-law phase. In contrast, strong tilts accelerate relaxation, leading to a single area-law phase. Moreover, due to the unique properties of the tilted chain, we examine the impact of the boundary on entanglement relaxation. This work not only deepens our understanding of the role of feedback in measurement-induced entanglement phase transitions but also enhances our comprehension of the non-Hermitian skin effect beyond the no-click limit.
@@ I Introduction
Depicting the evolution of quantum matter far from equilibrium with universal quantum dynamics might be notoriously challenging. However, considerable efforts to understand out-of-equilibrium dynamics in lattice models [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11], randum unitary circuits [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22], black hole information theory [23, 24, 25], and quantum field theory [26, 27, 28] reveal that universal patterns may emerge in the entanglement structures of many-body systems. The core issues in condensed matter physics today, such as quantum thermalization [29, 30, 31], quantum chaos [32, 33, 34], holography [28, 35], and information scrambling [36, 37, 38], all live in the shadow of these entanglement patterns evolution.
@@
In unitary entanglement relaxation dynamics (hereafter refer to entanglement relaxation), local perturbations spread entanglement within the Lieb-Robinson light cone [39], initially increasing entanglement entropy until it saturates. Growth patterns vary by system type: integrable systems show slower growth [40, 41, 42, 43, 44], non-integrable systems exhibit faster thermalized growth [45, 46], and many-body localized systems display logarithmic growth due to localization effects [47, 46, 48]. Interestingly, introducing weak measurements leads to a competition between unitary and non-unitary dynamics in the system, resulting in diverse entanglement relaxation patterns as well, thereby driving measurement-induced entanglement phase transitions (MIPT) [12, 13, 14, 49, 50, 51, 52, 53, 18].
@@
On the one hand, in the no-click limit [54], MIPT corresponds to purely non-Hermitian dynamics, where the post-selection mechanism of measurement-feedback effectively erases trajectories with quantum jumps. This phenomenon often directly reduces to an entanglement dynamics problem involving a Hatano-Nelson type [55, 56, 57] effective Hamiltonian, where entanglement relaxation is accompanied by a steady-state entanglement phase transition, eventually equilibrating to a non-Hermitian skin state [58, 59]. The entanglement relaxation and MIPT without quantum jumps have been extensively studied in free lattice models[60, 61, 11, 62, 63, 64], even including its interaction with localization potentials [60, 61, 65, 11]. However, implementing post-selection experimentally is notably difficult. On the other hand, entanglement relaxation and MIPT under quantum jumps [66] can also be described through stochastic dynamics in quantum many-body trajectories. In this context, if the effective Hamiltonian exhibits many-body skin pumping, quantum jump noise drives the entanglement steady state into a highly entangled configuration, rather than allowing it to settle into a low-entanglement non-Hermitian skin state. Consequently, the information scrambling from quantum jumps makes it challenging to observe the many-body skin effect in experimental settings.
However, substantial recent efforts, have been made to understand entanglement relaxation with quantum jumps under generalized monitoring [67], especially within the context of the skin effect, in lattice models and quantum circuits [68, 69, 67, 70, 71]. There, feedback operations suppress the dynamical noise caused by quantum jumps, allowing the system’s long-time steady-state to be dominated by the skin effect of the non-Hermitian effective Hamiltonian. Under the rescaled time ${\rm time}/L$, where $L$ is the system size, the feedback-induced skin effect drives non-monotonic bipartite entanglement growth. Even more surprisingly, this process reveals a novel dynamical transition that exhibits robustness against disordered-induced localization perturbations [70]. Nevertheless, whether this emergent entanglement dynamical transition induced by monitoring-feedback exists in disorder-free localized systems, particularly in systems with a Wannier-Stark (or tilted) potential [72], has not been studied before. The Wannier-Stark potential not only leads to hyperexponentially localized wave functions distinct from disorder in single-particle systems but also offers significant numerical and theoretical advantages, as it eliminates the need for disorder averaging. Additionally, with the rise of quantum simulation experimental platforms [73, 74], discussions on the non-equilibrium behavior of quantum many-body systems are no longer confined to Hilbert space, and disorder-free localization systems offer promising experimental prospects. A natural question that arises is: what is the fate of feedback-induced skin effects in the entanglement relaxation under the Wannier-Stark potential?
In this work, we analyze monitoring-feedback induced entanglement relaxation in a tilted free fermionic chain using the quantum trajectory method. Averaging over multiple Stochastic Schrödinger Equation (SSE) evolutions with quantum jumps on the time scale ${\rm time}/L$, we observe bipartite entanglement dynamics under a Wannier-Stark potential. Initially, entanglement enters a transient metastable phase, followed by long-time relaxation to a low-entanglement steady state, akin to the skin effect. With minor tilting, non-monotonic entanglement relaxation and dynamical transition persist. However, strong tilting appears to accelerate relaxation, potentially suppressing the entanglement transition. Additionally, we examine boundary conditions, highlighting insights into feedback-monitored entanglement dynamics and the non-Hermitian skin effect with quantum jumps. Our findings not only shed light on the entanglement relaxation dynamics of feedback-monitored quantum many-body systems under the competition of disorder-free localization but also contribute to a deeper understanding of the non-Hermitian skin effect with quantum jumps.
@@
This paper is organized as follows. Section II introduces the model along with the monitoring-feedback protocol. In Section III, we present the analysis of trajectory-averaged relaxations, focusing on the time evolution of the average density distribution. Section IV provides numerical evidence of the entanglement dynamical transitions under varying tilts. Additionally, we explore the quasiparticle interpretation of the bulk and edge dynamics in Section V. Finally, Section VI offers a summary and outlines future research directions. Additional numerical calculation data and discussions are included in the Appendix.
@@ II Setups
In this work, we consider a tilted free fermion chain of length $L$ and choose open boundary conditions (OBCs) unless otherwise specified. The system is described by the following Hamiltonian,
$\hat{H}=\sum_{l}^{L-1}{J(\hat{c}_{l}^{\dagger}\hat{c}_{l+1}+H.c.)}+\Delta\sum_{l}^{L}{l\hat{c}_{l}^{\dagger}\hat{c}_{l}},$ (1)
where $\hat{c}_{l}^{\dagger}$ ($\hat{c}_{l}$) is the fermion creation (annihilation) operator at site $l$, $J$ is the strength of nearest-neighbor hopping which is set to $1$ hereafter for convenience, and $\Delta$ is the strength of the tilted potential, i.e., the on-site potential increases linearly with the site index $l$ with the potential difference between the nearest-neighbor sites being $\Delta$. Introducing the continuous measurements, the evolution of the open quantum system is described by the Lindblad master equation which is a formidable challenge for numerical simulation. Thus, to simply this process, we instead consider the SSE [66, 75, 76, 77, 78, 79] corresponding to the trajectory evolution of a pure state $\left|\psi(t)\right\rangle$
@@
where $\hat{L}_{l}$ is a quantum jump operator induced by weak measurement, $\langle\hat{O}\rangle=\left\langle\psi(t)\right|\hat{O}\left|\psi(t)\right\rangle$ is the expectation value of operator $\hat{O}$, $dW_{l}$ is an independent and discrete Poisson random variable, which takes values from {0, 1} with mean value $\langle\hat{L}_{l}^{\dagger}\hat{L}_{l}\rangle\gamma dt$. $\hat{H}_{\text{eff}}$ is non-Hermitian effective Hamiltonian, $\hat{H}_{\text{eff}}=\hat{H}-i\gamma/{2}{\textstyle\sum_{l}^{L-1}{\hat{L}_{l}^{\dagger}\hat{L}_{l}}}$, where $\gamma$ is the monitoring rate.
Figure 1: (a) The influence of the continuous measurement with the feedback under the tilted potential. (b) OBCs, the dynamical phase diagram of the trajectory averaged bipartite entanglement entropy in the plane (${\rm time}/L$, $\Delta$) with $L=192$. The figure is primarily segmented by log-law region, area-law region, and no exact edge.
@@
where $e^{i\theta\hat{n}_{l+1}}$ is the feedback (see Appendix B for the feedback protocol considering the first site under PBCs), $\hat{n}_{l+1}$ is the particle number operator, and $\hat{d}^{\dagger}_{l}=(\hat{c}^{\dagger}_{l}-i\hat{c}^{\dagger}_{l+1})/\sqrt{2}$ is the creation operator of a right-moving quasimode. If $\theta\neq 0$, the feedback will convert the right-moving quasimode to the left-moving one, and we set $\theta=\pi$ without loss of generality. Based on the previous text, we now can write $\hat{H}_{\text{eff}}$ [see FIG. 1 (a)] as
@@
In the subsequent sections, we set $\gamma=0.5$, which corresponds to the case of the critical entanglement phase caused by the weak measurement.
Figure 2: The trajectory averaged density distribution evolution with respect to rescaled time for $L=128$. (a) OBC, $\Delta=0$; (b) OBC, $\Delta=0.6$; (c) PBC, $\Delta=0$; (d) PBC, $\Delta=0.6$.
@@
We choose the Neél state as the initial state, i.e., $\left|\psi(t=0)\right\rangle=\left|0101...01\right\rangle$. Since both $\hat{H}_{\text{eff}}$ and $\hat{L}_{l}$ are quadratic, we can utilize the Gaussian state to efficiently simulate each trajectory evolution of $\left|\psi(t)\right\rangle$. We note the observables are averaged 200-800 trajectories, represented by the variable with an overline (e.g., trajectory averaged density distribution $\overline{\langle\hat{n}_{l}\rangle}$). Moreover, the observables can be easily obtained from the correlation matrix $C_{ij}=\langle\hat{c}^{\dagger}_{i}\hat{c}_{j}\rangle$. $\langle\hat{n}_{l}\rangle$ are the diagonal elements of $C_{ij}$. For the trajectory averaged bipartite entanglement entropy $\overline{S}_{L/2}$  [80, 81, 82, 83, 84], we first choose the left half chain as the subsystem, then get sub correlation matrix $C_{ij}^{L/2}$ from $C_{ij}$ and lastly obtain ${S}_{L/2}$ based on all eigenvalues $\{\lambda_{i}\}$ of $C_{ij}^{L/2}$ with $S_{L/2}=-{\textstyle\sum_{i}{\lambda_{i}\log{(\lambda_{i})}}+(1-\lambda_{i})\log{(1-\lambda_{i})}}$. The trajectory-averaged velocity of particles, $\overline{v}$, as derived from the effective Hamiltonian $\hat{H}_{\text{eff}}$ and the correlation function $C_{ij}$, reflects the dynamics of entanglement relaxation [70]. See Appendix A for more details.
III Trajectory averaged relaxations
Figure 3: (a) and (c) show the time evolution of the trajectory averaged bipartite entanglement entropy $\overline{S}_{L/2}$ for different $L$ under OBCs and PBCs separately. The green solid lines represent the absence of the tilted potential ($\Delta=0$), meanwhile, the orange solid lines represent the presence of the tilted potential ($\Delta=0.6$). The colors ranging from light to dark correspond to $L=64,96,128,160,192$. In the zoomed inset, the dashed black highlights the transition point, which separates the log-law and area-law regions. Inset: $\overline{S}_{L/2}$ versus $\log{L}$ at ${\rm time}/L=$ $0.5$ (diamond), $1.8$ (square). (b) and (d) illustrate the maximum bipartite entanglement entropy $\overline{S}_{L/2}^{max}$ with respect to $\Delta$ for $L=64,96,128,160,192$ (from light to dark colors). Inset: $\overline{S}_{L/2}^{max}$ versus $\log{L}$ for different $\Delta$.
@@ III Trajectory averaged relaxations
We first investigate how the time-evolving average density distribution $\overline{\langle\hat{n}_{l}\rangle}$, behaves under different boundary conditions: open boundary conditions (OBCs) and periodic boundary conditions (PBCs). Additionally, we will examine the effects of the presence and absence of a tilted potential. In FIG. 2 (a), (b), and (d), the late-time state is characterized by a notable feature: particles predominantly occupy the left half of the chain, approaching what is commonly referred to as a domain wall state however without a distinct boundary. The phenomenon observed is known as the feedback-induced skin effect, as detailed in Ref. [67]. Specifically, under OBCs, a system under the tilted potential with a strength of $\Delta=0.6$ exhibits a prolonged residence in the non-equilibrium state compared to a system without the tilted potential ($\Delta=0$). This extended duration in the non-equilibrium state is likely attributable to the restriction of the tilted potential for the particle motion.
@@
FIG. 2 (c) stands out from the others in the series, as it demonstrates that $\overline{\langle\hat{n}_{l}\rangle}$ remains relatively uniform across all lattice sites for the majority of the observation period. This uniformity can be attributed to the absence of edges and the dynamics are determined by the bulk thus preventing the skin effect. Furthermore, when comparing FIG.  2 (c) and (d), despite both under PBCs, the system in (d) does not exhibit such uniform distribution. Conversely, as time evolves the relaxation process closely mirrors the behavior observed in (b) and late-time $\overline{\langle\hat{n}_{l}\rangle}$ is mainly distributed on the left of the chain. The feedback-induced skin effect still exists in this system despite under PBCs. The reason is that the presence of the tilted potential creates a significant disparity between the first and the last on-site potentials, which generates a pseudo edge. Therefore, the particles hardly move from the first to the last site. Under PBCs, the pseudo edge plays a role similar to that of the edge under OBCs, restricting the bulk dynamics of particles. The dashed line shown in FIG. 2 indicates a transition point within the relaxation process, that will be further discussed in the following section.
IV Entanglement relaxations and dynamical transition

### v2

# Quantum feedback induced entanglement relaxation and dynamical phase transition in monitored free fermion chains with Wannier-Stark ladder

@@ Abstract
In recent years, measurement-induced entanglement transitions (MIET) have attracted significant attention. However, the dynamical transition associated with the feedback-induced skin effect, which exhibits a wealth of intriguing phenomena, has not been fully understood. In this work, we investigate a dynamical phase transition in a tilted free-fermion chain under measurement-feedback protocols, emphasizing particle density and entanglement entropy dynamics. We reveal a feedback-induced skin effect, enhanced by the Wannier-Stark ladder potential, which creates localization at one boundary and generates an effective pseudo-edge under periodic conditions. The observables show a two-stage evolution: a rapid initial logarithmic growth followed by decay into an area-law steady state. Using finite-size scaling analysis, we pinpoint the critical transition and provide an intuitive physical picture linking it to feedback-driven suppression of quantum jump fluctuations. While the Wannier-Stark ladder delays relaxation without changing universality, strong gradients can directly induce area-law entanglement, bypassing the dynamical transition. This work provides a effective supplement to the dynamical transition proposed in Ref. [1]. It provides valuable references for linking the dynamical understanding of the role feedback plays in MIET.
@@ I Introduction
Depicting the evolution of quantum matter far from equilibrium with universal quantum dynamics might be notoriously challenging. However, considerable efforts to understand out-of-equilibrium dynamics in lattice models [2, 3, 4, 5, 6, 7, 8, 9, 10, 11], random unitary circuits [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22], black hole information theory [23, 24, 25], and quantum field theory [26, 27, 28] reveal that universal patterns may emerge in the entanglement structures of many-body systems. The core issues in condensed matter physics today, such as quantum thermalization [2, 29, 30], quantum chaos [31, 32, 33], holography [28, 34], and information scrambling [35, 36, 37], are all characterized by distinct entanglement patterns.
@@
In unitary entanglement relaxation dynamics (hereafter refer to entanglement relaxation), local perturbations spread entanglement within the Lieb-Robinson light cone [38], initially increasing entanglement entropy until it saturates. Growth patterns vary by system type: integrable systems exhibit slower growth [39, 40, 41, 42, 43], non-integrable systems thermalize and show faster growth [44, 3], and many-body localized systems display logarithmic growth due to localization effects [45, 3, 46]. Interestingly, introducing weak measurements leads to a competition between unitary and nonunitary dynamics in the system, resulting in diverse entanglement relaxation patterns as well, thereby driving measurement-induced entanglement transitions (MIET) [12, 13, 14, 47, 48, 49, 50, 51, 18].
@@
On the one hand, in the no-click limit [52], MIET corresponds to purely non-Hermitian dynamics, where the post-selection mechanism of measurement feedback effectively erases trajectories with quantum jumps. This phenomenon often directly reduces to an entanglement dynamics problem involving a Hatano-Nelson type [53, 54, 55] effective Hamiltonian, where entanglement relaxation is accompanied by a steady-state entanglement phase transition, eventually equilibrating to a non-Hermitian skin state [56, 57]. The relaxation of entanglement and MIET without quantum jumps have been extensively studied in free lattice models[58, 59, 11, 60, 61, 62], even including its interaction with localization potentials [58, 59, 63, 11]. However, implementing post-selection experimentally is notably difficult. On the other hand, entanglement relaxation and MIET under quantum jumps [64] can also be described through stochastic dynamics in quantum many-body trajectories. In this context, if the effective Hamiltonian exhibits many-body skin pumping, quantum jump noise drives the entanglement steady state into a highly entangled configuration, rather than allowing it to settle into a low-entanglement non-Hermitian skin state. Consequently, the information scrambling from quantum jumps makes it challenging to observe the many-body skin effect in experimental settings.
However, substantial recent efforts have been made to understand entanglement relaxation with quantum jumps under generalized monitoring [65, 66, 67, 68], especially within the context of the skin effect, in lattice models and quantum circuits [69, 70, 66, 1, 71]. There, feedback operations suppress the dynamical noise caused by quantum jumps, allowing the system’s long-time steady-state to be dominated by the skin effect of the non-Hermitian effective Hamiltonian. Under the rescaled time $t/L$, where $L$ is the system size, the feedback-induced skin effect drives nonmonotonic bipartite entanglement growth, which has been identified as a novel dynamical transition [1]. However, several questions remain open following Ref. [1]: Is this truly a dynamical phase transition? What is the underlying physical mechanism of this transition? Does this transition depend on the rescaling of $t/L$?
More interestingly, the tilt system realized by introducing a linear potential into the Hamiltonian has emerged as a new playground for exploring exotic quantum states. Instead of undergoing unbounded acceleration, as one might expect from classical intuition, particles in a Wannier-Stark ladder can exhibit Bloch oscillations [72, 73, 74] or become localized in Wannier-Stark states [75, 76, 77]. In many-body settings, tilts can lead to Stark many-body localization [46, 78, 79] and Hilbert space fragmentation [80, 81, 82], thereby hindering thermalization and giving rise to rich non-equilibrium phenomena. Notably, with the advent of increasingly clean quantum simulation platforms [79, 81, 82, 78], tilt systems come with a lower experimental overhead compared to other localization potentials. A natural question that arises is whether the aforementioned dynamical transition [1] could emerge in the presence of a Wannier-Stark ladder, potentially giving rise to even more intriguing nonequilibrium phenomena.
In this work, we investigate a dynamical phase transition in a tilted (Wannier-Stark laddered) free-fermion chain under monitoring-feedback protocols, with a focus on the dynamics of particle density profiles and entanglement entropy. We discover a feedback-induced skin pumping effect, where the presence of the Wannier-Stark ladder extends the evolution time required for the initial density profile to transition into the final domain-wall configuration. Under periodic boundary conditions, this potential generates an effective boundary or pseudo-edge, facilitating similar localization even in the absence of physical edges. Our analysis of entanglement entropy indicates a two-step evolution: a rapid initial increase following logarithmic scaling, subsequently transitioning into a gradual decay toward an area-law entangled steady state. Using multiple observables combined with finite-size scaling, we accurately pinpoint the critical transition separating these dynamical regimes. Furthermore, we offer an intuitive picture connecting this behavior to feedback-driven suppression of quantum jump fluctuations. Importantly, the Wannier-Stark ladder delays system relaxation but does not alter the dynamical phase transition’s universality class. Nevertheless, strong gradient potentials can independently lead the system directly into area-law entanglement, thus circumventing an explicit dynamical transition. This work effectively complement the unresolved issues regarding the mechanism and criticality of dynamical phase transitions highlighted in Ref. [1]. It provides valuable references for the possible connection between MIET and non-Hermitian dynamics.
@@
This paper is organized as follows. Section II introduces the model along with the monitoring-feedback protocol. In Section III, we present a comprehensive analysis of the dynamical relaxation behavior in our system. We begin by examining the trajectory-averaged relaxations, focusing on the time evolution of the average density distribution. Next, we provide numerical evidence for the entanglement dynamical transitions under varying tilts, and discuss the evolution of other relevant observables during the relaxation process. Finally, using finite-size scaling, we analyze the critical parameters of this dynamical phase transition, which further confirms the universal nature of the observed behavior. Section IV provides an explanation for this dynamical phase transition, summarizes the paper, and outlines future research directions. Additional numerical calculation data and discussions are included in the Appendix.
Figure 1: (a) The influence of the continuous measurement with the feedback under the Wannier-Stark ladder. (b) OBCs, the dynamical phase diagram of the trajectory-averaged bipartite entanglement entropy in the plane ($t/\tau$, $\Delta$) with $L=192$. The figure is primarily segmented by log-law region and area-law region.
@@ II Setups
II.1 Model
In this work, we consider a free fermion chain of length $L$ with Wannier-Stark ladder. The system is described by the following Hamiltonian,
$\hat{H}=\sum_{l}{J(\hat{c}_{l}^{\dagger}\hat{c}_{l+1}+H.c.)}+\Delta\sum_{l}{l\hat{c}_{l}^{\dagger}\hat{c}_{l}},$ (1)
where $\hat{c}_{l}^{\dagger}$ ($\hat{c}_{l}$) is the fermion creation (annihilation) operator at site $l$, $J$ is the strength of nearest-neighbor hopping which is set to $1$ hereafter for convenience, and $\Delta$ is the gradient of the Wannier-Stark ladder, i.e., the on-site potential increases linearly with the site index $l$ with the potential difference between the nearest-neighbor sites being $\Delta$. Introducing the continuous measurements, the evolution of the open quantum system is described by the Lindblad master equation, which is a formidable challenge for numerical simulation. Further, intrinsic property of the trajectory state is not visible in the properties of the average state.
Thus, to simplify this process, we instead consider the SSE [64, 83, 84, 85, 86, 87] corresponding to the trajectory evolution of a pure state $\left|\psi(t)\right\rangle$
@@ II.1 Model
where $\hat{L}_{l}$ is a quantum jump operator induced by weak measurement, $\langle\hat{O}\rangle=\left\langle\psi(t)\right|\hat{O}\left|\psi(t)\right\rangle$ is the expectation value of operator $\hat{O}$, $dW_{l}$ is an independent and discrete Poisson random variable, which takes values from {0, 1} with mean value $\langle\hat{L}_{l}^{\dagger}\hat{L}_{l}\rangle\gamma dt$. $\hat{H}_{\text{eff}}$ is non-Hermitian effective Hamiltonian, $\hat{H}_{\text{eff}}=\hat{H}-i\gamma/{2}{\textstyle\sum_{l}^{L-1}{\hat{d}_{l}^{\dagger}\hat{d}_{l}}}$, where $\gamma$ is the monitoring rate and $\hat{d}^{\dagger}_{l}=(\hat{c}^{\dagger}_{l}-i\hat{c}^{\dagger}_{l+1})/\sqrt{2}$ is the creation operator of a right-moving quasimode.
@@
where $e^{i\theta\hat{n}_{l+1}}$ is the feedback (see Appendix B for boundary effect discussions), $\hat{n}_{l+1}$ is the particle number operator. If $\theta\neq 0$, the feedback will convert the right-moving quasimode to the left-moving one, and we set $\theta=\pi$ without loss of generality (for more discussion on theta, please refer to Appendix D). Based on the previous text, we now can write $\hat{H}_{\text{eff}}$ [see FIG. 1 (a)] as
@@
In the subsequent sections, we set $\gamma=0.5$ unless otherwise specified, which corresponds to the case of the critical entanglement phase caused by the weak measurement [88].
We choose the Néel state as the initial state, i.e., $\left|\psi(t=0)\right\rangle=\left|0101...01\right\rangle$. Since both $\hat{H}_{\text{eff}}$ and $\hat{L}_{l}$ are quadratic, we can utilize the Gaussian state to efficiently simulate each trajectory evolution of $\left|\psi(t)\right\rangle$. We note the observables are averaged 200-800 trajectories, represented by the variable with an overline (e.g., trajectory-averaged density distribution $\overline{\langle\hat{n}_{l}\rangle}$). We describe the evolution of the observables using the rescaled time $t/\tau$, where $\tau$ is a dimensionless coefficient proportional to $L$.
II.2 Observables
@@
To characterize the dynamics and entanglement properties of the system, we consider several key observables derived from the fermionic correlation matrix. The observables can be easily obtained from the correlation matrix $C_{ij}=\langle\hat{c}^{\dagger}_{i}\hat{c}_{j}\rangle$. For instance, $\langle\hat{n}_{l}\rangle$ are the diagonal elements of $C$. For the trajectory-averaged bipartite entanglement entropy $\overline{S}_{L/2}$ [89, 90, 91, 92, 93], we first choose the left half chain as the subsystem, then get sub-correlation matrix $C_{ij}^{L/2}$ from $C$ and lastly obtain ${S}_{L/2}$ based on all eigenvalues $\{\lambda_{i}\}$ of $C_{ij}^{L/2}$ with $S_{L/2}=-{\textstyle\sum_{i}{[\lambda_{i}\log{(\lambda_{i})}}+(1-\lambda_{i})\log{(1-\lambda_{i})]}}$. The trajectory-averaged velocity of particles, $\overline{v}$, as derived from the effective Hamiltonian $\hat{H}_{\text{eff}}$ and the correlation function $C_{ij}$, reflects the dynamics of entanglement relaxation [1]. The trajectory-averaged mutual information $\overline{I}_{AB}$ indicates the connection between two different subsystems, and its definition is given as $I_{AB}=S_{A}+S_{B}-S_{AB}$, with $S_{A}$, $S_{B}$ and $S_{AB}$ representing the entanglement entropies of subsystems $A$, $B$ and their union $A\cup B$, respectively. $\overline{f}_{\text{skin}}$ quantifies the time-averaged probability of the evolving state being in the ideal many-body skin state with $f_{\text{skin}}=|\left\langle\psi_{\text{skin}}|\psi(t)\right\rangle|^{2}$. The Frobenius norm $\overline{\|D\|}_{F}$ quantifies the subsystem correlations upon bipartitioning the system. Here, $D$ is extracted from the correlation matrix $C$ by selecting the submatrix $D=C_{L/2:L,1:L/2}$, which encodes cross-partition correlations. See Appendix A for more numerical details.
III Numerical Results
III.1 Trajectory-averaged relaxations
@@
We first investigate how the time-evolving trajectory-averaged density distribution $\overline{\langle\hat{n}_{l}\rangle}$, behaves under different boundary conditions: open boundary conditions (OBCs) and periodic boundary conditions (PBCs). Additionally, we will examine the effects of the presence and absence of a Wannier-Stark ladder. In FIG. 2 (a), (b), and (d), the late-time state is characterized by a notable feature: particles predominantly occupy the left half of the chain, approaching what is commonly referred to as a domain wall state, however, without a distinct boundary (hereafter termed the skin-like state). The phenomenon observed is known as the feedback-induced skin effect, as detailed in Ref. [66]. Specifically, under OBCs, a system under the Wannier-Stark ladder with a strength of $\Delta=0.6$ exhibits a prolonged residence in the non-equilibrium state compared to a system without the Wannier-Stark ladder ($\Delta=0$). This extended non-equilibrium duration is likely due to the restriction imposed by the Wannier-Stark ladder on particle motion.
@@
FIG. 2 (c) stands out from the others in the series, as it demonstrates that $\overline{\langle\hat{n}_{l}\rangle}$ remains relatively uniform across all lattice sites for the majority of the observation period. This uniformity can be attributed to the absence of edges and the dynamics are determined by the bulk, thus preventing the skin effect. Furthermore, when comparing FIG.  2 (c) and (d), despite both under PBCs, the system in (d) does not exhibit such uniform distribution. Conversely, as time evolves the relaxation process closely mirrors the behavior observed in (b) and late-time $\overline{\langle\hat{n}_{l}\rangle}$ is mainly distributed on the left of the chain. The feedback-induced skin effect still exists in this system despite under PBCs. The reason is that the presence of the Wannier-Stark ladder creates a significant disparity between the first and the last on-site potentials, which generates a pseudo edge. Therefore, the particles hardly move from the first to the last site. Under PBCs, the pseudo edge plays a role similar to that of the edge under OBCs, restricting the bulk dynamics of particles. The dashed lines shown in FIG. 2 (a), (c) and (d) indicate a transition point within the relaxation process, that will be further discussed in the following section.
Figure 2: Panels (a)-(d) decipt the evolution of the trajectory-averaged density distribution for $L$ = $128$. (a) OBC, $\Delta$ = 00; (b) OBC, $\Delta$ = $0.6$; (c) PBC, $\Delta$ = 00; (d) PBC, $\Delta$ = $0.6$. Panels (e) and (h) show the evolution of the $\overline{S}_{L/2}$ for different $L$ under OBCs and PBCs. The green solid lines represent the absense of the Wannier-Stark ladder ($\Delta$ = 00), while the orange solid lines denote its presence ($\Delta$ = $0.6$). The colors rangeing from light to dark correspond to $L$ = $64$, $96$, $128$, $160$ and $192$. In the zoomed inset, we highlight the transition by the dashed black line. We also show the $\overline{S}_{L/2}$ versus $\log{L}$ at $t/\tau$ = $0.5$(diamond), $1.8$(square). Panels (f) and (i) illustrate the maximum bipartite entanglement entropy $\overline{S}^{\text{max}}_{L/2}$ with respect to $L$ = $64$, $96$, $128$, $160$ and $192$ (from light to dark colors). The insets show the $\overline{S}^{\text{max}}_{L/2}$ versus $\log{L}$ for different $\Delta$ under OBCs and PBCs separately. Panel (g) shows the evolution of $\overline{v}$ for different $\Delta$. Panel (j) shows the evolution of $\overline{v}$ for different $\gamma$ under $\Delta$ = 00. Both panels (g) and (h) have a system size of $64$.
III.2 Entanglement relaxations and dynamical phase transitions